find the sum of the infinite geometric series: 9 + 6 + 4 + ...
Answer by benazir.sj@gmail.com(109)

find the sum of the infinite geometric series: 9 + 6 + 4 + ...
sum of infinite = a/ (1-r)
common ratio = 2nd / 1st
therefore
common ratio= 6/9=2/3
a=9,r=2/3
sum =a/(1-r)
=9/(1-2/3)
=9/(3-2/3)
=9/1/3
=9*3/1
27